OpenStudy (anonymous):

can someone help me with coming up with a real world problem for cubic functions?

OpenStudy (anonymous):

The open box problem comes to mind

OpenStudy (anonymous):

how does that work ?

OpenStudy (anonymous):

it's usually some variation of having a sheet of metal or cardboard where you cut out a corner and then maximize the volume of the box

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you could do an initial value problem involving second integrals? with acceleration integrated to velocity and then to position?

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

lets say you have acceleration = x + 1, then velocity is the integral of acceleration = (1/2)(x^2) + x + c and position is the integral of velocity = (1/6)(x^3) + (1/2)x^2 + cx + b

OpenStudy (anonymous):

ok i think i get it